# Efficient Redundancy Reduction for Open-Vocabulary Semantic Segmentation

### v1

@@ Abstract
Open-vocabulary semantic segmentation (OVSS) is an open-world task that aims to assign each pixel within an image to a specific class defined by arbitrary text descriptions. Recent advancements in large-scale vision-language models have demonstrated their open-vocabulary understanding capabilities, significantly facilitating the development of OVSS. However, most existing methods suffer from either suboptimal performance or long latency. This study introduces ERR-Seg, a novel framework that effectively reduces redundancy to balance accuracy and efficiency. ERR-Seg incorporates a training-free Channel Reduction Module (CRM) that leverages prior knowledge from vision-language models like CLIP to identify the most relevant classes while discarding others. Moreover, it incorporates Efficient Semantic Context Fusion (ESCF) with spatial-level and class-level sequence reduction strategies. CRM and ESCF result in substantial memory and computational savings without compromising accuracy. Additionally, recognizing the significance of hierarchical semantics extracted from middle-layer features for closed-set semantic segmentation, ERR-Seg introduces the Hierarchical Semantic Module (HSM) to exploit hierarchical semantics in the context of OVSS. Compared to previous state-of-the-art methods under the ADE20K-847 setting, ERR-Seg achieves +$5.6\%$ mIoU improvement and reduces latency by $67.3\%$. The project page is available at https://lchen1019.github.io/ERR-Seg.
Index Terms:
Open-vocabulary semantic segmentation, vision-language model, redundancy reduction
I Introduction
@@
Semantic segmentation is a fundamental task in computer vision that aims to achieve pixel-level semantic understanding. Modern semantic segmentation methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] rely heavily on large-scale labeled datasets, but can only achieve understanding of limited predefined categories. In contrast, humans can understand scenes in an open-vocabulary manner, enabling the comprehension of arbitrary categories. The pursuit of open-vocabulary semantic segmentation (OVSS) is to replicate this intelligence that can perform unbounded pixel-level semantic understanding. Encouragingly, through natural language supervision, recent large-scale visual-language pre-training models (VLMs) exhibit open-vocabulary comprehension ability. However, existing VLMs receive image-level supervision, which is inadequate for the pixel-level representation required by semantic segmentation. This representation granularity gap impedes the transfer of open-world knowledge from VLMs to OVSS.
@@
Figure 1: Performance vs. latency on ADE20K-847. Compared with ZegFormer [11], OVSeg [12], DeOP [13], SAN [14], SED [15] and CAT-Seg [16], ERR-Seg achieves a new state-of-the-art mIoU of $16.9\%$ with lower latency.
@@
For this issue, recent efforts can be categorized into mask-based and cost-based methods. Mask-based methods [11, 17, 12, 18, 19, 14, 20, 21, 22, 23] model OVSS as the combination of mask generation and open-vocabulary mask classification, these approaches require region-level representation, which is much closer to the image-level representation. However, mask-based methods depend on a region proposal network (RPN), which is typically trained on a specific dataset, lacking the region proposal capability in the open-vocabulary context. To address this issue, some attempts [24, 25, 23, 26] employ the SAM [27] as a universal RPN. Nevertheless, as SAM tends to over-segment an instance into sub-patches, these methods have not made significant progress yet. In contrast, cost-based methods [16, 15] directly generate coarse-grained cost maps by calculating the cosine similarity between dense pixel-level embeddings and text embeddings. These methods effectively bridge the representation granularity gap and eliminate the requirement of RPN. As a result, the methods in this family have shown superior performance compared to mask-based methods.
@@
Although cost-based methods have effectively addressed the challenge of representation granularity, there are still some issues to be further considered: (1) Redundancy: It is evident that real-world categories are limitless, while those depicted in an image are always limited, leading to numerous redundant classes. OVSS tasks struggle with this redundancy. On the one hand, it leads to increased latency due to the redundant channel dimension of the cost maps. On the other hand, it poses a challenge for the attention mechanism to model class-level sequences containing numerous redundant classes. (2) Hierarchical Semantic: Existing closed-set semantic segmentation methods [28, 2] typically extract hierarchical semantic maps from middle-layer features and aggregate them to derive the final semantic maps. In contrast, cost-based methods only integrate semantic information from pixel-text cost maps, neglecting the semantic details in middle-layer features.
@@
To this end, we propose ERR-Seg, a novel cost-based framework for OVSS that jointly considers efficiency and accuracy. To address the issue (1), ERR-Seg presents the Channel Reduction Module (CRM) to select the most relevant categories while discarding others based on the category existence probability implied in the cost maps. By reducing the channel dimensions, the attention mechanism can better capture spatial-level long-range dependencies, as depicted in Fig. 2. This reduction also significantly decreases latency. Furthermore, ERR-Seg introduces Efficient Semantic Context Fusion (ESCF) with Spatial-level Sequence Reduction (SSR) and Class-level Sequence Reduction (CSR), which reduces computational workload by $49.3\%$ without compromising performance. To address the issue (2), ERR-Seg incorporates the Hierarchical Semantic Module (HSM) to transform middle-layer features into semantic space to compute cost maps in parallel. These hierarchical cost maps are then merged by using a pyramid structure to aggregate more middle-layer semantic details. As shown in Fig. 1, ERR-Seg establishes a new state-of-the-art in terms of efficiency and accuracy.
@@
  • We propose a training-free Channel Reduction Module (CRM) to eliminate redundant classes. By customizing a small-scale vocabulary list for each image, our proposed ERR-Seg can handle an extensive number of classes while significantly reducing computational complexity.
  • We quantitatively and qualitatively explore the impact of redundant classes on spatial-level and class-level contextual information modeling in the context of OVSS.
  • We present the Hierarchical Semantic Module (HSM) to generate hierarchical cost maps, enabling the aggregation of semantic details from middle-layer features.
@@
  • We introduce an efficient framework for OVSS, which achieves superior performance on multiple datasets and significantly accelerates the inference process.
II Related Work
II-A Closed-set Semantic Segmentation
From the perspective of modeling semantic segmentation tasks, technical methods can be divided into two categories: (1) FCN-based methods [29] consider the semantic segmentation task as a dense pixel-by-pixel prediction task. Various FCN-based variants have been developed to incorporate more contextual information, such as dilated convolution [1, 7], feature pyramid [30, 28], pyramid pooling module [31, 28], object contextual representation [8], attention mechanisms [5, 32]. Additionally, some methods [6, 33, 2, 34, 9, 10] introduce transformer-based architectures to semantic segmentation. (2) Mask-based methods [35, 36, 3, 4] treat the segmentation task as a combination of mask generation and mask classification. These methods aim to unify various segmentation tasks (such as instance segmentation, semantic segmentation, and panoptic segmentation) and have demonstrated their effectiveness in the context of closed-set semantic segmentation.
Figure 2: Visual comparison of attention maps. With our proposed channel reduction module (CRM), the attention mechanism can more effectively capture spatial-level long-range dependencies in the context of OVSS.
Figure 3: Overall architecture of ERR-Seg. We first utilize the Hierarchical Semantic Module to generate dense pixel-text cost maps. The training-free Channel Reduction Module is then used to eliminate redundant classes. The Efficient Semantic Context Fusion further enhances the cost maps, improving spatial and class-level contextual information. Finally, the Gradual Upsampling Decoder restores the high-rank information of cost maps by incorporating image details from the middle-layer features $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$ of CLIP’s visual encoder.
II-B Vision-Language Models
Visual-language models (VLMs) aim to extend open-world comprehension to the visual domain by jointly training visual and language modalities. Early VLMs [37, 38, 39, 40] were trained on small-scale datasets through multiple multi-modal tasks. In contrast, CLIP [41] collects large-scale image-text pairs and uses image-text contrastive learning to align the visual and language modality. To further enhance the quality and efficiency of image-text pre-training, ALIGN [42] collected a much larger dataset of image-text pairs without removing noisy data, SigLIP [43] introduces pairwise sigmoid loss to eliminate the requirement of a global view of the pairwise similarities for normalization. Beyond image-text contrast learning, BLIP-2 [44] and LLaVA [45] introduce an additional language model to process aligned images and text in a unified framework for more complex reasoning. In this study, we further explore the use of prior knowledge from VLMs to better support the OVSS task.
II-C Open-Vocabulary Semantic Segmentation
Earlier methods [46, 47, 48] for OVSS attempt to establish a joint embedding space connecting image pixels and a predefined vocabulary list. However, these methods struggle with arbitrary class recognition due to their limited vocabulary size. Fortunately, vision-language models (VLM) [41, 42, 49] pre-trained on extensive web data exhibit strong capabilities in open-vocabulary recognition, leading recent research focused on adapting these models for OVSS. Based on the method of introducing semantic information from VLM, current OVSS methods can be categorized into two main groups: mask-based methods and cost-based methods.
For mask-based methods, an intuitive two-stage approach is first proposed. Initially, an additional RPN is adopted to propose region candidates, followed by the application of VLMs for open-vocabulary recognition. ZegFormer [11] and SimBaseline [17] directly adopt this two-stage method, leading to suboptimal performance and significant computational overload. To improve the performance, ODISE [19] utilizes a text-to-image diffusion model as RPN to generate high-quality masks, and OVSeg [12] fine-tunes CLIP on region-text pairs for improved open-vocabulary recognition. To further reduce the computational overload, single-stage methods have been introduced to avoid multiple forward propagations. Among these methods, MaskCLIP [50] and SAN [14] adopt attention bias, while FC-CLIP [20] employs mask pooling. To further enhance the performance of single-state methods, EBSeg [21] alleviates overfitting to training classes through image embedding balancing, and MROVSeg [22] introduces a multi-resolution training framework to tackle the problem that VLMs are typically pre-trained with low-resolution images.
For cost-based methods, CAT-Seg [16] aggregates the pixel-text cost maps at both spatial-level and class-level at one resolution. It also emphasizes the importance of fine-tuning the text encoder of CLIP. In contrast, SED [15] gradually aggregates the cost maps at different resolutions. However, cost-based methods have a drawback compared to mask-based methods: they involve operators with quadratic complexity in relation to the vocabulary size. This means that when the vocabulary is large, it can lead to significant computational and storage overhead. Our proposed ERR-Seg is also cost-based, but we introduce a training-free mechanism to remove redundant classes, thereby tailoring a small-scale vocabulary list for each image.
III Methodology
III-A Overall Architecture
The overview of our proposed ERR-Seg is shown in Fig. 3. Given an image $I\in\mathbb{R}^{3\times H\times W}$ and an arbitrary set of classes $\mathcal{C}=\{C_{1},...,C_{K}\}$, the OVSS task aims to assign a specific class $C_{i}$ to each pixel in $I$. Here, $K$ is the number of classes. Our framework comprises four stages: (1) Encoding the image $I$ to derive the middle-layer features $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$, as well as dense image embeddings $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$. Encoding the classes $\mathcal{C}$ to generate text embeddings of each class $\mathcal{T}$. (2) Generating the pixel-text cost maps by using $\mathcal{F}_{3}^{s}$, $\mathcal{F}_{4}^{s}$ and $\mathcal{T}$, and removing redundant classes to customize small-scale cost maps for each image. (3) Enhancing the cost maps by incorporating more spatial-level and class-level contextual information. (4) Leveraging the high-rank textual information from $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$, to generate a more detailed high-resolution segmentation result.
III-B Hierarchical Semantic Module
Typically, cost-based methods introduce semantic information by calculating pixel-text cost maps using only the last layer output, which ignores semantic details from middle-layer features. To address this, we propose the Hierarchical Semantic Module (HSM) to extract hierarchical semantic details from the middle-layer features of open-vocabulary visual encoders. HSM calculates pixel-text cost maps on the middle-layer features in parallel and merges these hierarchical cost maps through a feature pyramid strategy.
Initially, the HSM first encodes $I$ to derive the middle-layer features $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$, and dense image-text aligned embeddings $\mathcal{F}_{4}^{s}$. The relationship between $\mathcal{F}_{3}^{v}$ and $\mathcal{F}_{4}^{s}$ is modeled as follows:
$\mathcal{F}_{4}^{s}$ $=g\left(f\left(conv_{s=2}(\mathcal{F}_{3}^{v})\right)\right),$ (1)
where $f$ represents stage 4 of CLIP’s visual encoder, $g$ is the linear projection head, and $conv_{s=2}$ denotes convolution with a stride of 2 for downsampling. Additionally, the HSM generates extra image-text aligned embeddings from $\mathcal{F}_{3}^{v}$:
$\mathcal{F}_{3}^{s}=g\left(f\left(conv_{s=1}(\mathcal{F}_{3}^{v})\right)\right),$ (2)
where $conv_{s=1}$ denotes convolution with a stride of 1, using the same initial values as $conv_{s=2}$. $conv_{s=1}$ prevents downsampling, thereby keeping the size of $\mathcal{F}_{3}^{s}$ the same as $\mathcal{F}_{3}^{v}$.
Figure 4: Pipeline of our proposed CRM mechanism. CRM involves a training-free scoring function to assign scores to each class, retaining the top-$P$ classes while eliminating other redundant classes.
Subsequently, HSM computes the hierarchical cost maps. Given a vocabulary list $\mathcal{C}=\{C_{1},...,C_{K}\}$, a set of sentences $\mathcal{G}=\{G_{1},...,G_{K}\}$ is first generated, where $G_{i}$ consists of $C_{i}$ and a well-crafted prompt (e.g., “a photo of $\{C_{i}\}$ in the scene”). $\mathcal{G}$ is then fed into the CLIP’s text encoder to get text embeddings for each class, denoted as $\mathcal{T}=\{T_{1},...,T_{K}\}$. The pixel-text cost maps are derived by calculating the cosine similarity between each pixel and each class:
where $(i,j)$ denotes the 2D spatial positions, and $t$ represents the index of text embeddings. The $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$ are used to derive pixel-text cost maps $S_{3}$ and $S_{4}$ individually.
To aggregate the hierarchical cost maps, we employ a feature pyramid strategy. More specifically,
$S=conv\left(S_{3}\right)+up\left(conv\left(S_{4}\right)\right),$ (4)
where $S_{3}$ and $S_{4}$ are first projected into the same dimension by $conv$, then $S_{4}$ is upsampled to align with the size of $S_{3}$ via $up$, and subsequently add them to yield final cost maps $S\in\mathbb{R}^{K\times C\times H_{c}\times W_{c}}$.
III-C Channel Reduction Module
The Channel Reduction Module (CRM) is proposed to eliminate redundant classes, based on the observation that the number of classes within an image is limited. As illustrated in Table I, images annotated with $847$ classes have an average of only $10.2$ categories, indicating that there are $98.8\%$ redundant classes. Even images annotated with 20 classes have more than $90.0\%$ redundant classes. The impact of redundant classes will be discussed later in this section. CRM introduces a training-free channel selection function $\Phi$, which leverages the prior knowledge in original cost maps $S\in\mathbb{R}^{K\times C\times H_{c}\times W_{c}}$ to convert $S$ into reduced cost maps $S_{r}=\Phi\left(S\right)$, where $S_{r}\in\mathbb{R}^{K\times C\times H_{c}\times W_{c}}$ and $K>P$. $P$ is an empirical value and will be discussed later.
TABLE I: Analysis of image category count.
@@ III-C Channel Reduction Module
Redundancy $92.5\%$ $91.9\%$ $94.3\%$ $98.6\%$ $98.8\%$

### v2

@@ Abstract
Open-vocabulary semantic segmentation (OVSS) is an open-world task that aims to assign each pixel within an image to a specific class defined by arbitrary text descriptions. While large-scale vision-language models have shown remarkable open-vocabulary capabilities, their image-level pretraining limits effectiveness on pixel-wise dense prediction tasks like OVSS. Recent cost-based methods narrow this granularity gap by constructing pixel-text cost maps and refining them via cost aggregation mechanisms. Despite achieving promising performance, these approaches suffer from high computational costs and long inference latency. In this paper, we identify two major sources of redundancy in the cost-based OVSS framework: redundant information introduced during cost maps construction and inefficient sequence modeling in cost aggregation. To address these issues, we propose ERR-Seg, an efficient architecture that incorporates Redundancy-Reduced Hierarchical Cost maps (RRHC) and Redundancy-Reduced Cost Aggregation (RRCA). Specifically, RRHC reduces redundant class channels by customizing a compact class vocabulary for each image and integrates hierarchical cost maps to enrich semantic representation. RRCA alleviates computational burden by performing both spatial-level and class-level sequence reduction before aggregation. Overall, ERR-Seg results in a lightweight structure for OVSS, characterized by substantial memory and computational savings without compromising accuracy. Compared to previous state-of-the-art methods on the ADE20K-847 benchmark, ERR-Seg improves performance by $5.6\%$ while achieving a 3.1× speedup. The project page is available at https://lchen1019.github.io/ERR-Seg.
keywords:
[…]
1 Introduction
@@
Semantic segmentation is a fundamental task in computer vision that aims to achieve pixel-level semantic understanding. Modern semantic segmentation methods [chen2017rethinking, xie2021segformer, cheng2022masked, jain2023oneformer, fu2019dual, strudel2021segmenter, chen2018encoder, yuan2020object, chen2022vision, xia2024vit] rely heavily on large-scale labeled datasets, but can only achieve understanding of limited predefined categories. In contrast, humans can understand scenes in an open-vocabulary manner, enabling the comprehension of arbitrary categories. The pursuit of open-vocabulary semantic segmentation (OVSS) is to replicate this intelligence that can perform unbounded pixel-level semantic understanding. Encouragingly, through natural language supervision, recent large-scale visual-language pre-training models (VLMs) exhibit open-vocabulary comprehension ability. However, existing VLMs receive image-level supervision, which is inadequate for the pixel-level representation required by semantic segmentation. This representation granularity gap impedes the transfer of open-world knowledge from VLMs to OVSS.
@@
Figure 1: Performance vs. latency on ADE20K-847. Compared with ZegFormer [ding2022decoupling], OVSeg [liang2023open], DeOP [han2023open], SAN [xu2023side], SED [xie2024sed] and CAT-Seg [cho2024cat], ERR-Seg achieves a new state-of-the-art with lower latency.
@@
For this issue, recent efforts can be categorized into mask-based and cost-based methods. Mask-based methods [ding2022decoupling, xu2022simple, liang2023open, dong2023maskclip, xu2023open, xu2023side, yu2024convolutions, shan2024open, zhu2024mrovseg, chen2024sam] conceptualize OVSS as the combination of mask generation and open-vocabulary mask classification. These methods necessitate region-level representation, which is much closer to the image-level representation. However, mask-based methods depend on a region proposal network (RPN) to generate masks, which is typically trained on a specific dataset, lacking the region proposal capability in the open-vocabulary context. To address this issue, several studies [chen2023semantic, ren2024grounded, chen2024sam, wang2024sam] employ SAM [kirillov2023segment] as a universal RPN. Nevertheless, as SAM tends to over-segment an instance into sub-patches, these methods have not made significant progress yet. In contrast, cost-based methods [cho2024cat, xie2024sed] directly generate coarse-grained cost maps by calculating the cosine similarity between dense pixel-level embeddings and text embeddings. These methods effectively bridge the representation granularity gap and eliminate the requirement of RPN. As a result, the methods in this family have shown superior performance compared to mask-based methods.
@@
Although cost-based methods have effectively addressed the challenge of representation granularity, they introduce cost aggregation over pixel-text cost maps, incurring significant computational overhead. Moreover, this overhead grows quadratically with the number of categories, in contrast to the linear growth of mask-based methods. We attribute this issue to two inherent design redundancies in the cost-based methods: (1) Redundancy in Cost Maps: It is evident that real-world categories are limitless, while those depicted in an image are always limited, leading to numerous redundant classes. As illustrated in Table 1, images annotated with $847$ classes have an average of only $10.2$ categories, indicating that there are $98.8\%$ redundant classes. On the one hand, it leads to increased latency due to the redundant class dimension of the cost maps. On the other hand, it poses a challenge for the attention mechanism to model class-level sequences containing numerous redundant classes. (2) Redundancy in Cost Aggregation: In closed-set semantic segmentation [fu2019dual, yuan2020object, xie2021segformer], the fixed class vocabulary enables attention mechanisms to jointly model spatial-class and class-level contextual information with class-informative channels. In contrast, the cost-based OVSS methods require class-agnostic channels due to dynamic class vocabulary, leading to the spatial-level and class-level decoupled attention mechanism in cost aggregation. While effective, it incurs significant computational and memory overhead. Additionally, due to the decoupling, the attention mechanism’s objective is relatively straightforward, potentially containing redundant computations that can be reduced.
@@
To this end, we propose an efficient framework in view of redundancy reduction for OVSS, jointly considering the efficiency and accuracy. To alleviate the redundancy (1), ERR-Seg presents the Redundancy-Reduced Hierarchical Cost maps (RRHC). It first selects the most relevant categories while discarding the others based on the category existence probability implied in the pixel-text cost maps. Additionally, it transforms middle-layer features from CLIP’s image encoder into semantic space to compute hierarchical cost maps in parallel, which are then merged to aggregate more middle-layer semantic details. By reducing the channel dimensions, the attention mechanism can better capture spatial-level long-range dependencies (as depicted in Fig. 2), while also significantly decreasing latency. More comprehensive analyses are provided in Sec. 3.4. To alleviate the redundancy (2), ERR-Seg introduces Redundancy-Reduced Cost Aggregation (RRCA), which incorporates spatial-level and class-level sequence reduction before cost aggregation. This results in a $49.3\%$ reduction in computational workload in cost aggregation without compromising performance. As demonstrated in Fig. 1, ERR-Seg establishes a new state-of-the-art in terms of efficiency and accuracy.
@@
  • We introduce an efficient cost-based OVSS framework in view of redundancy reduction, which achieves superior performance on multiple datasets and significantly accelerates the inference process.
  • We introduce Redundancy-Reduced Hierarchical Cost maps (RRHC), characterized by dynamic redundant class reduction for each image and hierarchical cost maps extracted from middle-layer features.
  • We present the Redundancy-Reduced Cost Aggregation (RRCA) with spatial-level and class-level sequence reduction, substantially accelerating the aggregation process without compromising accuracy.
  • We quantitatively and qualitatively investigate the impact of redundancy reduction on modeling spatial-level and class-level contextual information, as well as its effect on computational complexity.
Table 1: Analysis of image category count across five semantic segmentation datasets. It reveals that both PAS-20 (with 20 categories) and A-847 (with 847 categories) contain more than $90\%$ redundant classes.
@@
Redundancy $92.5\%$ $91.9\%$ $94.3\%$ $6\% 98.89$$8\%$
Figure 2: Visual comparison of attention maps during cost aggregation. (a) Using our proposed cost maps (RRHC) with redundancy reduction (48 classes) and (b) using the original full cost maps (847 classes). The results show that the attention mechanism more effectively captures spatial-level long-range dependencies in our proposed RRHC.
2 Related Work
2.1 Closed-set Semantic Segmentation
From the perspective of modeling semantic segmentation tasks, technical methods can be divided into two categories: (1) Pixel-based methods consider the semantic segmentation task as a dense pixel-by-pixel prediction task. FCN [long2015fully] serves as a fundamental pixel-based approach by substituting the fully connected layer in conventional CNNs with a convolutional layer to enable dense category predictions. Subsequently, various pixel-based variants have been developed to incorporate more contextual information, such as dilated convolution [chen2017rethinking, chen2018encoder], feature pyramid [kirillov2019panoptic, xiao2018unified], pyramid pooling module [zhao2017pyramid, xiao2018unified], object contextual representation [yuan2020object], attention mechanisms [fu2019dual, huang2019ccnet], pixel contrast [wang2021exploring], prototype view [zhou2022rethinking] and hierarchical semantic [li2022deep, li2023logicseg]. Additionally, some methods [strudel2021segmenter, zheng2021rethinking, xie2021segformer, gu2022multi, chen2022vision, xia2024vit, liang2023clustseg] introduce transformer-based architectures to semantic segmentation. (2) Mask-based methods [cheng2021per, zhang2021k, cheng2022masked, jain2023oneformer] treat the segmentation task as a combination of mask generation and mask classification. These methods aim to unify various segmentation tasks (such as instance segmentation, semantic segmentation, and panoptic segmentation) and have demonstrated their effectiveness in the context of closed-set semantic segmentation.
As a fundamental task in computer vision, closed-set semantic segmentation has achieved remarkable success in the past decade, particularly in specific domains such as remote sensing [ma2024sam, geng2023dual], medical segmentation [zhou2024uncertainty] and autonomous driving scenarios [wu2024s, chang2023multi]. However, this approach relies heavily on large-scale labeled datasets, which require significant human effort to create. Moreover, it is limited to recognizing only a predefined set of categories. This study focuses on OVSS, which can understand arbitrary vocabulary lists, allowing it to address more complex and dynamic scenarios.
Figure 3: Overall architecture of ERR-Seg. Initially, redundancy-reduced hierarchical cost maps are generated by extracting cost maps from middle-layer features and eliminating class redundancy. Subsequently, the sequence length is reduced before cost aggregation to speed up the computation. Finally, the upsampling decoder restores the high-rank information of cost maps by incorporating image details from the middle-layer features of CLIP’s visual encoder.
2.2 Vision-Language Models
Visual-language models (VLMs) aim to extend open-world comprehension to the visual domain by jointly training visual and language modalities. Early VLMs [chen2019uniter, su2019vl, lu2019vilbert, li2020oscar] were trained on small-scale datasets through multiple multi-modal tasks. In contrast, CLIP [radford2021learning] collects large-scale image-text pairs and uses image-text contrastive learning to align the visual and language modality. To further enhance the quality and efficiency of image-text pre-training, ALIGN [jia2021scaling] collected a much larger dataset of image-text pairs without removing noisy data, SigLIP [zhai2023sigmoid] introduces pairwise sigmoid loss to eliminate the requirement of a global view of the pairwise similarities for normalization. Beyond image-text contrast learning, BLIP-2 [li2023blip] and LLaVA [liu2024visual] introduce an additional language model to process aligned images and text in a unified framework for more complex reasoning.
VLMs have been successfully utilized in various multi-modal downstream tasks such as text-to-image generation [ramesh2022hierarchical, zhang2023adding, wang2024prior], referring segmentation [wang2022cris, shang2022cross], cross-modal retrieval [cheng2022semantic, li2024ckdh] and open-vocabulary understanding [liu2024grounding, yu2024convolutions, xie2024sed]. In this study, we better support the application of VLMs to OVSS by enhancing their ability to generate hierarchical semantic maps and leveraging the prior knowledge from VLMs for redundancy reduction.
2.3 Open-Vocabulary Semantic Segmentation
Earlier methods [bucher2019zero, xian2019semantic, zhao2017open] for OVSS attempt to establish a joint embedding space connecting image pixels and a predefined vocabulary list. However, these methods struggle with arbitrary class recognition due to their limited vocabulary size. Fortunately, vision-language models (VLM) [radford2021learning, jia2021scaling, li2022blip] pre-trained on extensive web data exhibit strong capabilities in open-vocabulary recognition, leading recent research focused on adapting these models for OVSS. Based on the method of introducing semantic information from VLM, current OVSS methods can be categorized into two main groups: mask-based methods and cost-based methods.
For mask-based methods, an intuitive two-stage approach is first proposed. Initially, an additional RPN is adopted to propose region candidates, followed by the application of VLMs for open-vocabulary recognition. ZegFormer [ding2022decoupling] and SimBaseline [xu2022simple] directly adopt this two-stage method, leading to suboptimal performance and significant computational overload. To improve the performance, ODISE [xu2023open] utilizes a text-to-image diffusion model as RPN to generate high-quality masks, and OVSeg [liang2023open] fine-tunes CLIP on region-text pairs for improved open-vocabulary recognition. To further reduce the computational overload, single-stage methods have been introduced to avoid multiple forward propagations. Among these methods, MaskCLIP [ding2022open] and SAN [xu2023side] adopt attention bias, while FC-CLIP [yu2024convolutions] employs mask pooling. Recent works have introduced various techniques to enhance performance. EBSeg [shan2024open] alleviates overfitting to seen training classes via image embedding balancing. MROVSeg [zhu2024mrovseg] tackles the resolution discrepancy by introducing a multi-resolution training framework. SCAN [liu2024scan] and MAFT+ [jiao2024maftp] enhance the fusion of visual and textual features through semantic guidance and a collaborative optimization mechanism, respectively. More recently, OVSNet [liu2025stepping] further avoids overfitting by fusing heterogeneous features and expanding the training space.
For cost-based methods, CAT-Seg [cho2024cat] aggregates the pixel-text cost maps at both the spatial-level and class-level at one resolution. It also emphasizes the importance of fine-tuning the text encoder of CLIP. In contrast, SED [xie2024sed] gradually aggregates the cost maps at different resolutions. However, cost-based methods have a drawback compared to mask-based methods: they involve operators with quadratic complexity in relation to the vocabulary size. This means that when the vocabulary is large, it can lead to significant computational and storage overhead. Our proposed ERR-Seg is also cost-based, but we propose to reduce redundancy to alleviate this issue.
3 Methodology
Figure 3 presents an overview of our proposed ERR-Seg. To address the inherent redundancy within the cost-based framework for OVSS, we propose eliminating redundancies in both cost maps and cost aggregation. This section details our redundancy reduction approach for cost maps in Sec. 3.1, followed by the method for reducing redundancy in cost aggregation in Sec. 3.2. Additionally, Sec. 3.3 provides a brief introduction to the upsampling encoder adopted in the cost-based framework. Finally, Sec. 3.4 presents a qualitative and quantitative analysis of the impact of redundancy.
3.1 Redundancy-Reduced Hierarchical Cost Maps
In this section, we present the methodology for constructing our proposed Redundancy-Reduced Hierarchical Cost maps (RRHC). It diverges from previous cost maps in two aspects: (1) Hierarchical Cost Maps: previous methods typically utilize only the last layer output from the CLIP’s image encoder, whereas RRHC also incorporates the features from intermediate layers to build hierarchical cost maps. (2) Redundant Class Reduction: we eliminate redundant classes to derive more concise cost maps customized for each image.
3.1.1 Hierarchical Cost Maps
Typically, cost-based methods introduce semantic information by calculating pixel-text cost maps using only the last layer output, which ignores semantic details from middle-layer features. To address this, we extract semantic details from the middle-layer features from the CLIP’s image encoders to build hierarchical cost maps.
Initially, $I$ is first encode to derive the middle-layer features $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$, and dense image-text aligned embeddings $\mathcal{F}_{4}^{s}$. The relationship between $\mathcal{F}_{3}^{v}$ and $\mathcal{F}_{4}^{s}$ is modeled as follows:
$\mathcal{F}_{4}^{s}$ $=f_{\mathrm{head}}\left(g\left(conv_{s=2}(\mathcal{F}_{3}^{v})\right)\right),$ (1)
where $g$ represents stage 4 of CLIP’s visual encoder, $f_{\mathrm{head}}$ is the linear projection head, and $conv_{s=2}$ denotes convolution with a stride of 2 for downsampling. Additionally, we generate extra image-text aligned embeddings from middle-layer features $\mathcal{F}_{3}^{v}$, as follows:
$\mathcal{F}_{3}^{s}=f_{\mathrm{head}}\left(g\left(conv_{s=1}(\mathcal{F}_{3}^{v})\right)\right),$ (2)
where $conv_{s=1}$ denotes convolution with a stride of 1, using the same initial values as $conv_{s=2}$. $conv_{s=1}$ prevents downsampling, thereby keeping the size of $\mathcal{F}_{3}^{s}$ the same as $\mathcal{F}_{3}^{v}$. Notably, we do not construct cost maps from $\mathcal{F}_{1}^{v}$ or $\mathcal{F}_{2}^{v}$, as these high-resolution features primarily capture high-frequency details. Therefore, incorporating them during the upsampling stage is sufficient.
Subsequently, by giving $\mathcal{F}_{3}^{v}\in\mathbb{R}^{C\times\frac{H}{16}\times\frac{W}{16}}$ and $\mathcal{F}_{4}^{v}\in\mathbb{R}^{C\times\frac{H}{32}\times\frac{W}{32}}$, we computes the hierarchical cost maps. Given a vocabulary list $\mathcal{C}=\{C_{1},...,C_{K}\}$, a set of sentences $\mathcal{G}=\{G_{1},...,G_{K}\}$ is first generated, where $G_{i}$ consists of $C_{i}$ and a well-crafted prompt (e.g., “a photo of $\{C_{i}\}$ in the scene"). $\mathcal{G}$ is then fed into the CLIP’s text encoder to get text embeddings for each class, denoted as $\mathcal{T}=\{T_{1},...,T_{K}\}$. The pixel-text cost maps are derived by calculating the cosine similarity between each pixel and each class:
where $(i,j)$ denotes the 2D spatial positions, and $t$ represents the index of text embeddings, $\mathcal{F}^{s}=\mathcal{F}_{3}^{s}$ or $\mathcal{F}_{4}^{s}$. The $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$ are used to derive the pixel-text cost maps $S_{3}\in\mathbb{R}^{K\times\frac{H}{16}\times\frac{W}{16}}$ and $S_{4}\in\mathbb{R}^{K\times\frac{H}{32}\times\frac{W}{32}}$ individually, combining into our proposed hierarchical cost maps.
Figure 4: Pipeline of our proposed redundant class reduction mechanism. It involves a training-free scoring function to assign scores to each class, retaining the top-$P$ classes while eliminating other redundant classes.
3.1.2 Redundant Class Reduction
Based on the observation that the number of classes within an image is limited, we introduce a training-free class selection function $\Phi$, which leverages the prior knowledge in the original cost maps to reduce redundant classes, and determines the top-$P$ classes. Here, $P\leq K$ and $P$ is an empirical value. The impact of $P$ on training and inference will be discussed later.